OpenStudy (diyadiya):

The given function \[f(x)= \log (x+ \sqrt{x^2+1})\] is an even function ? odd function ? or a periodic function ?

OpenStudy (anonymous):

do you know how to test if function is even odd?

OpenStudy (diyadiya):

Yeah f(x)= f(-x) Even f(x)=f(x) Odd

OpenStudy (anonymous):

if you just put x=1 and x=-1 and check with calculator results are same just one positive and other negative so it's odd, but idk how to show with x :/

OpenStudy (anonymous):

\[f(-x)=\log(-x+\sqrt{x^2+1})\]

OpenStudy (anonymous):

@ishaan94 knows :D

OpenStudy (diyadiya):

Ishaan's not here !

OpenStudy (anonymous):

you need to show that \[\log(-x+\sqrt{x^2+1})=-\log(x+\sqrt{x^2+1})\] but i don't know how :/

OpenStudy (diyadiya):

okay :)

OpenStudy (anonymous):

None :/

OpenStudy (apoorvk):

any even function is that, where f(x)=f(-x), that is, basically, the function takes same values for any particular 'x' as well as '-x' so, the function's graph is symmetrical about the Y-axis. something like this... (considering any arbitrary function)  so you may have guessed what you have to do in here.. tell me if i need to explain more.